OpenStudy (anonymous):

what is x^2-16/x-4 simplified

OpenStudy (cwrw238):

= 3(x^2 - x - 20) - can you factor whats in the parentheses?

Parth (parthkohli):

Hmm Idk what cwrw told, but here's what you should do. \(x^2 - 1 6 = (x + 4)(x - 4)\). The numerator is factored.

Parth (parthkohli):

\( \color{Black}{\Rightarrow\Large{(x + 4)(x - 4) \over (x - 4)} }\)

Parth (parthkohli):

If you want to know more about how \(x^2 - 16 = (x + 4)(x - 4)\), see my tutorial here: http://openstudy.com/updates/4fd62be0e4b04bec7f17374c